Skip to main content
A-LevelMathematicsProbability and statistics9709/61 · Oct/Nov 2024 · Q714 Marks

The number of accidents per year on a certain road has the distribution Po(λ). In the past the value of λ was 3.3. Recently, a new speed limit was imposed and the council wishes to test whether the value of λ has decreased. The council notes the total number, X, of accidents during two randomly chosen years after the speed limit was introduced and it carries out a test at the 5% significance level.

  1. (a)Calculate the probability of a Type I error.[4]
  2. (b)Given that X = 2, carry out the test.[3]
  3. (c)The council decides to carry out another similar test at the 5% significance level using the same hypotheses and two different randomly chosen years. Given that the true value of λ is 0.6, calculate the probability of a Type II error.[3]
  4. (d)Using λ = 0.6 and a suitable approximating distribution, find the probability that there will be more than 10 accidents in 30 years.[4]

Mark scheme

Marking points from the official mark scheme (Oct/Nov 2024). The mark or mark type is shown in brackets.

  1. [B1] For two years, λ = 3.3 * 2 = 6.6
  2. [M1] P(X < 2) = e⁻⁶.⁶(1 + 6.6 + 6.6²/2!) [= 0.0400] [< 0.05 ] or e⁻⁶.⁶(1 + 6.6 + 21.78) or 0.001360 + 0.008978 + 0.02963
  3. [B1] P(X ≤ 3) = e⁻⁶.⁶(1 + 6.6 + 6.6²/2 + 6.6³/3!) or 0.0400 + e⁻⁶.⁶ 6.6³/3! = 0.105 [ > 0.05 ]
  4. [A1] P(Type I error) = 0.0400 (3 sf)
  5. [B1] H₀: λ = 6.6, H₁: λ < 6.6
  6. [M1] [P(X ≤ 2) = 0.0400], 0.04 < 0.05
  7. [A1] [Reject H₀] There is evidence to suggest that mean number of accidents has decreased
  8. [M1] P(X > 2) attempted, with any λ
  9. [M1] P(X > 2) = 1 - e⁻¹.²(1 + 1.2 + 1.2²/2!) or = 1 - e⁻¹.²(1 + 1.2 +0.72) or = 1- (0.3012 +0.3614 +0.2169 )
  10. [A1] 0.121 (3 sf) or 0.120
  11. [B1] For 30 years, λ = 0.6 * 30 = 18. N(18, 18) seen or implied
  12. [M1] P(Y > 10.5) = (10.5 - 18) / √18 [= -1.768]
  13. [M1] P(X > '-1.768') = Φ('1.768')
  14. [A1] = 0.961 or 0.962 (3 sf)

Write your answer and get it marked

On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.

Sign up free →

About this question

Topic
Probability and statistics, A-Level Mathematics (syllabus 9709). Structured question, 14 marks.
Source
A-Level Mathematics, Oct/Nov 2024 Paper 6 Variant 1, question 7.

Questions and mark schemes are from past examination papers and remain the copyright of the exam board. How Oracle Prep builds its content

Related Mathematics questions

© 2026 Oracle Prep